00Bastian Gruberx

 00  x
 00  x

00Professional x 00LinkedIn: linkedin.com/in/bastiangruberx 00Book: rustwebdevelopment.com x 00E-Mail: gruberbastian@me.com x 00Website: bastiangruber.ca x 00Code x 00GitHub: github.com/gruberb x 00Git: code.novanexus.ca/gruberbx

SUMMARY

Staff Engineer at Mozilla, shipping Rust, C++, and Python to hundreds of millions of Firefox users across Desktop, Android, and iOS. Promoted for leading cross-team projects and getting complex, multi-platform features shipped. 12+ years of backend and systems work across distributed services, Kubernetes-based infrastructure, and production operations; published Rust author and long-time technical writer.

SKILLS

Languages: Rust, C++, Python, TypeScript, Go Systems & Backend: async Rust (Axum, Actix), Python/FastAPI, Node.js, FFI/UniFFI, gRPC, REST

Architecture: distributed systems, P2P protocols, system design, cross-platform (Desktop/Android/iOS)

Infrastructure & Observability: AWS, GCP, Docker, Kubernetes, CI/CD, Grafana, performance profiling

PROFESSIONAL EXPERIENCE

Staff Software Engineer Feb 2026 - Present

Mozilla, Remote Promoted for cross-team technical leadership and delivering multi-platform privacy features.

Leading the Image Service project: replaced 3 fragile, duplicated favicon systems with a single GCS-backed API serving 4+ internal teams. Extended an existing Python/Airflow pipeline instead of building a new service, reducing maintenance burden across the board

Implemented Oblivious HTTP (OHTTP) in application-services (shared Rust component cross-compiled to Swift/Kotlin via UniFFI) and wrote the C++ integration for Firefox Desktop. Now live on Firefox iOS, Android, and Desktop, covering millions of users

Maintaining the Firefox Push notification service (Rust backend on Kubernetes, 50k req/s, sub-200ms latency). Reduced memory footprint, added circuit breakers, built Grafana dashboards for observability Senior Software Engineer Jun 2024 - Jan 2026

Mozilla, Remote

Core maintainer of Merino, the Kubernetes-deployed Python/FastAPI service powering Firefox Suggest. All search suggestions in Firefox route through Merino, which enriches and serves them in a privacy-preserving way

Code reviewer for UniFFI, Mozilla’s multi-language bindings generator (Rust to Kotlin, Swift, Python, JavaScript)

Contributing directly to Firefox Desktop (mozilla-central): shipping Rust and C++ code

Led “Rusty Fridays” sessions (6–10 engineers per session), building internal Rust knowledge across teams

Technical Editor, “From C to Rust” Feb 2026 - Present

Manning Publications (Contract, Part-time) Reviewing technical accuracy and pedagogical quality for Manning’s C-to-Rust migration guide.

Senior Systems Engineer, Team Lead Apr 2023 - Jun 2024

Toposware (acquired by Polygon Labs), Remote Core team of 4 building a peer-to-peer distributed network protocol from scratch in Rust, with Kubernetes-based deployment and benchmarking environments. Our core implementation of the double echo protocol and the Ethereum bridge became part of the Polygon stack after acquisition.

Achieved 10x message throughput per node by profiling the actual bottleneck: optimized batch processing and serialization paths after discovering nodes were compute-bound on crypto operations, not network-limited

Built the EVM bridge connecting external Ethereum-compatible chains to the core network

Led the internal benchmarking initiative that changed where the team focused optimization effort

Senior Software Engineer Aug 2022 - Feb 2023

Centrifuge, Remote Protocol team building DeFi infrastructure on Substrate/Polkadot in Rust.

Implemented on-chain fund pool mutations, the core feature enabling live capital allocation adjustments in the lending protocol Replaced manual documentation with CI-driven auto-generation from Rust source, eliminating persistent drift between code and public docs

Author, “Rust Web Development” Oct 2020 - Dec 2022

Manning Publications Wrote and published an end-to-end guide covering async Rust, web frameworks, and production deployment.

Sr. Solutions Architect (promoted from SA) Jun 2020 - Jul 2022

Twilio Inc., Remote

Technical lead for Fortune 500 professional services engagements: diagnosed latency issues, redesigned messaging architectures, delivered implementation roadmaps sized to each client’s team and resources

Created reusable engagement playbooks adopted across the EMEA Professional Services team

Systems Engineer, Core Backend Aug 2019 - May 2020

Kraken Digital Asset Exchange, Remote Part of the first Rust team rewriting Kraken’s core infrastructure from PHP.

Rewrote core banking services and API gateway in Rust, replacing legacy PHP services. The new stack handled significantly higher throughput with a smaller hardware footprint

Shipped the first version of the automated KYC verification flow, replacing manual processing right as COVID drove a massive spike in new customer sign-ups

Earlier Career 2013 - 2019

Sr. Software Engineer (freelance from 2016) across enterprise clients (OSRAM, Kxrber Digital, GIATA, Avenga) and startups (Dalia Research, blogfoster, smartB) in Berlin.

Founded “Rust & Tell Berlin” meetup, grew it into the largest Rust meetup globally. Coordinated with companies for venue sponsorship, organized speakers, photographers, and videographers. Gave talks and created a welcoming space for the Rust community

Designed IoT authentication infrastructure at OSRAM for thousands of connected devices (Node.js, Go, Lua)

First engineering hire at Kxrber Digital: built real-time sensor data pipelines and cloud service deployments on AWS (Lambda, ECS, Route53, Kubernetes)

Decomposed GIATA’s monolithic PHP system into 6 Docker-based services, taught the existing team Scrum and JavaScript

Contributed to the Rust compiler and standard library during a self-directed sabbatical (2019)

Technical Writer 2008 - 2020

Freelance (Contract) 13 years writing for technical publications including c’t, iX, Macwelt (IDG/Heise), and tech blogs (LogRocket, stadt-bremerhaven.de, mobilegeeks.de). Also authored “OS X fxr Einsteiger” (Bookboon, 2012).

EDUCATION

Bachelor of Science, Computer Science Mar 2014

TH Ingolstadt. Scholarship from Allianz Deutschland AG (fully funded dual-study program, 2010–2014)